OpenStudy (anonymous):

what is slope intercept form? i mean when they are wanting me to express it in slope intercept form from looking at a graph is that different than saying (3,-4)?

OpenStudy (anonymous):

Slope intercept form is the form of y=mx + b where the slope is m, and the y-intercept is b. Slope can be calculated as follows: \[m=\frac{ y _{2}-y _{1}}{ x _{2}-x _{1} }\]The y-intercept is the point in which the graph hits the vertical axis of the y-axis. So that means if you have a y-intercept at point b, then it is located at point (0,b) for whatever value b is.

OpenStudy (anonymous):

so if on my graph the x axis is marked at -4 and the y axis is marked at 1 how would i write that out then

OpenStudy (anonymous):

Ah, well then they give you two points, so you are able to find the point. The graph on the x-axis is marked at -4, so that's an x-intercept, which can be written as the following point (-4,0). Same with the y-axis, which is marked at 1 as a y-intercept, written as (0,1). So just calculate the slope, then put it in the y=mx+b form.

OpenStudy (anonymous):

so would it be y=-4x+1???

OpenStudy (phi):

I think you did the slope upside down

OpenStudy (anonymous):

I believe that you have it reversed. The top is the Y values, and the bottom is the X values, when calculating the slope.

OpenStudy (anonymous):

what lol

OpenStudy (anonymous):

so it is y=1x+-4

OpenStudy (phi):

the y-intercept is 1 (your first equation)

OpenStudy (anonymous):

Try calculating the slope using the actual slope formula I provided earlier.

OpenStudy (phi):

the slope is the difference between the y values divided by difference of x values

OpenStudy (anonymous):

im confused ugh

OpenStudy (phi):

your 2 pts are (-4,0) and (0,1)

OpenStudy (anonymous):

correct

OpenStudy (phi):

the 2nd number is the y

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

(-4,1)

OpenStudy (phi):

so 0-1 is the y difference -4 -0 is the x difference

OpenStudy (phi):

you divide to get a singe fraction -1/-4 = 1/4

OpenStudy (anonymous):

oh ok that makes more sense

OpenStudy (phi):

y= (1/4)x +1 if x is 0 you get y= (1/4)*0+1 and y=0+1=1 so (0,1) is on the line if x= -4 y= (1/4)*-4 +1 = -4/4 + 1 = -1+1 =0 so (-4,0) is on the line looks good

OpenStudy (anonymous):

awesome thx